OpenStudy (anonymous):

Which percent is equivalent to 3? 300% 0.03% 3%

OpenStudy (jdoe0001):

\(\bf {\color{red}{ a}}\% =\cfrac{{\color{red}{ a}}}{100}\)

OpenStudy (jdoe0001):

ohh.ahemm 3.. ok ...

OpenStudy (anonymous):

Help

OpenStudy (jdoe0001):

\(\bf {\color{red}{ a}}\% =\cfrac{{\color{red}{ a}}}{100}\qquad \qquad {\color{red}{ a}}\%\times 100 = {\color{red}{ a}}\\ \quad \\ {\color{red}{ 3}}\% =\cfrac{{\color{red}{ 3}}}{100}\qquad \qquad {\color{red}{ 3}}\%\times 100 = {\color{red}{ 3}}\)

OpenStudy (anonymous):

is the answer 3% ?

OpenStudy (anonymous):

this is hard D:

OpenStudy (jdoe0001):

well..ahemm .. 3 * 100 =?

OpenStudy (jdoe0001):

\(\bf {\color{red}{ 3}}\% =\cfrac{{\color{red}{ 3}}}{100}\qquad \qquad {\color{red}{ 3}}\%\times 100 = {\color{red}{ 3}} \implies 300\%=3\)

OpenStudy (anonymous):

O.O soo its 300% :D

OpenStudy (anonymous):

i clicked finished and i got 100% THANK YEW PPLZ

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

my like is saved!!